Find the values of all the trigonometric functions from the given information.
cos t = −3/5
, π ≤ t ≤ 3π/2

Answer by lwsshak3(11628)

Find the values of all the trigonometric functions from the given information.
cos t = −3/5
, π ≤ t ≤ 3π/2
***
reference angle t is in quadrant III where sin<0, cos<0
working with (3-4-5) reference right triangle in quadrant III
cos t=-3/5
sin t=-4/5
tan t=-4/-3=4/3
sec t=-5/3
csc t=-5/4
cot t=3/4